# The Potential Impact of Primordial Black Holes on Exoplanet Systems

### v2

@@ 1 Introduction
In many cases, the interaction of a planet-star system with its environment can be reduced to a gravitational three-body problem. One such situation is when the third body makes a passing flyby, which perturbs the planetary system. Such flybys exchange energy with the planet-star system and may perturb the orbit of the planet. Other scenarios include collision between the transient flyby with the planet or the dissociation of the planet-star system; for a review, see e.g. Cuello, et al. (2023). There are a variety of free-floating bodies which can potentially intrude on a stellar system, including free-floating planets, planetesimals, main sequence stars, brown dwarf stars, astrophysical black holes and binaries, as well as hypothetical populations of primordial black holes (PBH) or dark matter microhalos.
There have been many studies of how the eccentricity and semi-major axis distributions for exoplanetary systems may be impacted by encounters with stars. Good starting points in the existing literature on close encounters between stellar systems and intruding bodies include Hills (1975) and Li & Adams (2015). Moreover, Spurzem et al. (2009) explored analytical estimates of the changes in planetary orbital eccentricity and semi-major axis due to adiabatic and impulsive encounters with stellar bodies. Our aim is to extend the study of close encounters to consider more exotic hypothetical astrophysical bodies, focusing on PBH. One reason that populations of exotic objects are interesting is due to the fact that their typical mass scale, spatial distribution, and velocity dispersions may be very different to conventional objects such as stars and rouge planets. PBH are black holes which form not through stellar collapse, but rather from extreme overdensities in the early universe (Zel’dovich & Novikov, 1967; Hawking, 1971), as such PBH can have masses well below one Solar Mass ($M_{\odot}$).
@@
Figure 1: Plot made through “exoplanets.eu” with the selection “mass:mjup $>$ 0.1 AND ”confirmed” in planet__\__status”.
@@ 2 Simulating Close Encounters
We initialize REBOUND for a flyby object on a hyperbolic trajectory given the flyby’s mass, velocity at infinity, and impact parameter. Adding an object to a REBOUND simulation with a fully determined initial orbit requires the object’s semi-major axis, eccentricity, inclination $i$, longitude of the ascending node $\Omega$, argument of perihelion $\omega$, and true anomaly $\nu$. Figure 2 illustrates the trajectories traced out under our simulation for a single example run.
@@ 4.1 Dynamics across 10 Orders of Magnitude
Figure 4: Results of Gala simulations. The star is assumed to have a circular orbit around the galactic center at a distance $d$; we consider $d=$2,3,4, and 5 kpc. The left panel shows the average number of close encounters as a function of the distance of the planet-star system from the galactic centre. The centre and right panels show histograms of the distance between the star and the intruder, and their relative velocities. The histograms show combined results for an equal number of runs at each value of $d$.
@@
For part A, to estimate the number of PBH that enter the neighborhood of a given star—which we take to be a sphere of radius $r_{c}=30$ pc—we implement numerical galactic dynamics simulations using the public Gala code (Price-Whelan, 2017). This Astropy-affiliated Python package numerically integrates the trajectories of stars and other astrophysical objects based on a given mass model. These computations are detailed in Section 4.2.
Then, for part B, we analyze how many of these PBH that come within 30 pc (as defined by our 4.2 simulation) lie on trajectories that would lead to a close encounter with a planet-star system. The distribution of distances is informed by the closest approaches found in our Gala simulations. Thus, we estimate the likelihood that a given star would undergo a PBH flyby that could non-negligibly perturb its planetary system. We take the radius at which non-negligible flybys may occur to be 15 AU of the star for Jupiter-like planets, and 90 AU for Neptune-like planets. Any PBH that comes within this distance we record and model in the next step. This calculation is detailed in Section 4.3.
@@
This three-step subdivision is necessary to make the problem computationally tractable. Note that the scale of the galaxy is $\sim 1$ Mpc, while stellar systems are of order $10^{-4}$ parsec (this is the Neptune-Sun distance), thus the problem spans 10 orders of magnitude in distances. Compartmentalizing the problem into three units essentially tracks the relevant dynamics first at the galactic scale using Gala, then at the parsec scale via solid angle scaling calculations, and, finally, at the interplanetary scale $\lesssim 50$ AU using REBOUND. Figure 3 shows the flow diagram of our approach.
@@ 4.2 Simulating Galactic Dynamics
$f_{\rm PBH}\equiv\frac{\Omega_{\rm PBH}}{\Omega_{\rm DM}}\sim\frac{N_{\rm PBH}M_{\rm PBH}}{10^{12}M_{\odot}}\sim 10^{-7}\left(\frac{M_{\rm PBH}}{10^{-1}M_{\odot}}\right).$ (8)
@@
For substellar mass PBH, observational constraints on PBH (e.g. microlensing) are satisfied for sub-percent level abundances $f_{\rm PBH}\lesssim 10^{-2}$ (Carr & Kuhnel, 2020). Additionally, if dark matter can annihilate to Standard Model states, then this can lead to limits, in particular, from observations of extragalactic $\gamma$-rays (Ackermann et al, 2015). For $f_{\rm PBH}\sim 10^{-7}$ such indirect detection bounds are typically satisfied (Lacki & Beacom, 2010; Boucenna et al, 2017; Ginés, Witte, & Mena, 2022; Kadota & Tashiro, 2022; Chanda, Scholtz, & Unwin, 2022; Chanda, Mukherjee, & Unwin, 2025).
@@ 4.3 Probability of Close Encounter
For example, the probability that a flyby at initial distance 30 pc from the star passes within 15 AU of the star is roughly $p\sim p_{r_{c}}\sim 10^{-12}$. Figure 5 shows the probability that a PBH at initial distance $d=20,25,30$ pc will pass within a given distance of closest approach ($\alpha$) of a star.
@@ 5 Distributions of Exoplanet Orbits Following Encounters
Observe that the perturbation strength decreases exponentially with respect to increases in the flyby velocity until the threshold value $R_{v}\approx 10$, whereafter the perturbation strength is insensitive to further increases in the flyby velocity. For a Sun-Jupiter system, an intruder of mass $M_{I}=10^{-2}M_{\odot}$ and velocity $|v_{0}|\approx 130$ km/s can lead to percent-level changes in planetary eccentricities ($\Delta e\sim 0.01$).
@@
Finally, we next turn to part C of Section 4 and examine how PBH flybys may alter the statistical distribution of exoplanet orbits. For each set of initial orbital parameters, we implement a Monte Carlo sampling technique to randomly choose the parameters of the flyby. For star-planet systems, the impact parameter $b$ of the flyby dictates the distance of closest approach and is drawn from a random distribution such that $b_{i}=r_{i}\times\sqrt{U[0,1]}$, where $r_{i}$ is the radius of critical interaction distance, and $U[0,1]$ is a random variable from a uniform distribution between 0 and 1. For star-Jupiter systems $r_{J}=15~{}{\rm AU}$, whereas for star-Neptune systems $r_{N}=90$ AU. The angular parameters $\Omega$, $\varpi$, and inclination $i$ of the flyby are all uniformly drawn from $[-\pi,\pi]$. Moreover, each initial setup is ran for $10$ samples, with each run independently sampling the angular orbital parameters.
Our findings are presented in Figure 9, which shows the distribution of changes to the eccentricity for Jupiter and Neptune-like planets assuming a PBH intruder with mass $10^{-1}M_{\odot}$. Changes to Jupiter’s eccentricity are strongly peaked at very small values, whereas Neptune has a broader distribution. We attribute these differences to the fact that Jupiter is twenty times more massive than Neptune.
@@ 6 Limitations
[…]
While single planet systems are thought to be common (this conclusion may be biased by observational capabilities), multi-planet systems are far from atypical. It has been estimated that multi-planet systems comprise around a quarter of all Kepler planets (Sandford et al., 2019). Kepler-90 (a Sun-like star 2.8 light years away) has eight planets (Shallue & Vanderburg, 2018). Even for single planet systems the planetary mass can range from $\sim 10^{-2}M_{\odot}$ (i.e. 10 $M_{J}$) [^6] At masses over $0.013M_{\odot}$ thermonuclear fusion to deuterium is typically possible. Although exoplanet.eu includes entries up to $0.08M_{\odot}$, the bodies above $0.013M_{\odot}$ are likely Brown Dwarfs., for instance ROXs 42Bb (Currie, Burrows, & Daemgen, 2014), down to $\sim 10^{-7}M_{\odot}$ being the mass of both Mercury and PSR B1257+12A (Konacki & Wolszczan, 2003).
Refer to caption Refer to caption
Figure 9: The distribution of the resulting eccentricities of Jupiter (left)/Neptune (right) planet-star systems after a close encounter with a flyby PBH of mass $10^{-1}M_{\odot}$ and velocity $200$ km/s (taking for Jupiter $b_{J}\in 15\sqrt{U[0,1]}$ and for Neptune $b_{N}\in 90\sqrt{U[0,1]}$).
@@ 7 Concluding Remarks
In this work we have utilysed the public package Gala to quantify the expected the number of PBH flybys for a given abundance and have shown that even with a small abundance ($f_{\rm PBH}\sim 10^{-7}$) a significant number of stars would have undergone a close encounter with a PBH. We then implemented a Monte Carlo study using a modified version of the public code Rebound to quantify how significantly the orbits would be perturbed. We leave for future work the interesting and feasible prospect of multiple PBH encounters on a system, and the prospect of a stellar encounter followed by PBH encounters. We would also like to examine more complicated stellar systems beyond the simplified Jupiter-Sun type system studied here.
@@
While we do not expect to be able to use exoplanet observations to place constraints in the near future, this work outlines the general principles of how one might use a future precision catalogue of exoplanets to discover or constrain populations of PBH. Finally, we note that if one could identify a class of clean planet-star systems that always formed with well-defined orbital parameters, these could provide an ideal testing ground for such dynamical constraints on PBH. Intriguingly, hot Jupiters seem to typically be found in single-planet systems (Wright et al., 2009; Steffen & Agol, 2005; Konacki et al., 2012), although exceptions have been observed; for instance, WASP-47b (Becker et al., 2015).
@@ Acknowledgements
JU is supported by NSF grant PHY-2209998 and thanks the Berkeley Center for Theoretical Physics, Queen’s College, Oxford, and Rudolf Peierls Centre for Theoretical Physics, of the University of Oxford for their hospitality during this work. LH thanks the MIT-PRIMES program. We are grateful to Hanno Rein and Jakub Scholtz for helpful discussions.
1 $10^{-5}$ 0.34
2 $10^{-10}$ 0.11
3 $10^{-15}$ 0.03
[…]
@@ Appendix B Multiple PBH Encounters
$n_{\rm PBH}(8~{}{\rm kpc})\sim 10^{-8}{\rm pc}^{-3}\left(\frac{f_{\rm PBH}}{10^{-7}}\right)\left(\frac{0.1M_{\odot}}{M_{\rm PBH}}\right)~{}.$ (11)
@@
$\Gamma_{\bullet}(8~{}{\rm kpc})t_{g}\simeq 10^{-5}\left(\frac{f_{\rm PBH}}{10^{-7}}\right)\left(\frac{0.1M_{\odot}}{M_{\rm PBH}}\right)~{}.$ (12)

### v3

@@ 1 Introduction
In many cases, the interaction of a planet-star system with its environment can be reduced to a gravitational three-body problem. One such situation is when the third body makes a passing flyby, which perturbs the planetary system. Such flybys exchange energy with the planet-star system and may perturb the orbit of the planet. Other scenarios include collision between the transient flyby with the planet or the dissociation of the planet-star system; for a review, see e.g. [Cuello, et al. (2023)]. There are a variety of free-floating bodies which can potentially intrude on a stellar system, including free-floating planets, planetesimals, main sequence stars, brown dwarf stars, astrophysical black holes and binaries, as well as hypothetical populations of primordial black holes (PBH) or dark matter microhalos.
There have been many studies of how the eccentricity and semi-major axis distributions for exoplanetary systems may be impacted by encounters with stars. Good starting points in the existing literature on close encounters between stellar systems and intruding bodies include [Hills (1975)] and [Li & Adams (2015)]. Moreover, [Spurzem et al. (2009)] explored analytical estimates of the changes in planetary orbital eccentricity and semi-major axis due to adiabatic and impulsive encounters with stellar bodies. Our aim is to extend the study of close encounters to consider more exotic hypothetical astrophysical bodies, focusing on PBH. One reason that populations of exotic objects are interesting is due to the fact that their typical mass scale, spatial distribution, and velocity dispersions may be very different to conventional objects such as stars and rouge planets. PBH are black holes which form not through stellar collapse, but rather from extreme overdensities in the early universe (Zel’dovich & Novikov, 1967; Hawking, 1971), as such PBH can have masses well below one Solar Mass ($M_{\odot}$).
@@
Figure 1: Plot made through “exoplanets.eu” with the selection “mass:mjup >> 0.1 AND ”confirmed” in planet_\_status”.
@@ 2 Simulating Close Encounters
We initialize REBOUND for a flyby object on a hyperbolic trajectory given the flyby’s mass, velocity at infinity, and impact parameter. Adding an object to a REBOUND simulation with a fully determined initial orbit requires the object’s semi-major axis, eccentricity, inclination $i$, longitude of the ascending node $\Omega$, argument of perihelion $\omega$, and true anomaly $\nu$. Figure 2 illustrates the trajectories traced out under our simulation for a single run.
@@ 4.1 Dynamics across 10 Orders of Magnitude
Figure 4: Results of Gala simulations. The star is assumed to have a circular orbit around the galactic center at a distance $d$; we consider d=d=2,3,4, and 5 kpc. The left panel shows the average number of close encounters as a function of the distance of the planet-star system from the galactic centre. The centre and right panels show histograms of the distance between the star and the intruder, and their relative velocities. The histograms show combined results for an equal number of runs at each value of $d$.
@@
For part A, to estimate the number of PBH that enter the neighborhood of a given star—which we take to be a sphere of radius $r_{c}=30$ pc—we implement numerical galactic dynamics simulations using the public Gala code (Price-Whelan, 2017). This Astropy-affiliated Python package numerically integrates the trajectories of stars and other astrophysical objects based on a given mass model. The computations are detailed in Section 4.2.
Then, for part B, we analyze how many of these PBH that come within 30 pc (as defined by our simulation) lie on trajectories that would lead to a close encounter with a planet-star system. The distribution of distances is informed by the closest approaches found in our Gala simulations. Thus, we estimate the likelihood that a given star would undergo a PBH flyby that could non-negligibly perturb its planetary system. We take the radius at which non-negligible flybys may occur to be 15 AU of the star for Jupiter-like planets, and 90 AU for Neptune-like planets. Any PBH that comes within this distance we record and model in the next step. This calculation is detailed in Section 4.3.
@@
This three-step subdivision is necessary to make the problem tractable. The scale of the galaxy is $\sim 1$ Mpc, while stellar systems are of order $10^{-4}$ parsec (this is the Neptune-Sun distance), thus the problem spans 10 orders of magnitude in distances. Compartmentalizing the problem into three units tracks the relevant dynamics first at the galactic scale using Gala, then at the parsec scale via solid angle scaling calculations, and, finally, at the interplanetary scale $\lesssim 50$ AU using REBOUND. Figure 3 shows the flow diagram of our approach.
@@ 4.2 Simulating Galactic Dynamics
[…]
@@
For substellar mass PBH, observational constraints on PBH (e.g. microlensing) are satisfied for sub-percent level abundances $f_{\rm PBH}\lesssim 10^{-2}$ (Carr & Kuhnel, 2020). Additionally, if dark matter can annihilate to Standard Model states, then this can lead to limits, in particular, from observations of extragalactic $\gamma$-rays (Ackermann et al, 2015). For $f_{\rm PBH}\sim 10^{-7}$ such indirect detection bounds exclude thermal (“WIMP”) dark matter with velocity-independent annihilation cross sections (Adamek et al, 2019; Lacki & Beacom, 2010; Boucenna et al, 2017; Ginés, Witte, & Mena, 2022; Chanda, Scholtz, & Unwin, 2022), however, other dark matter particles remain viable. Examples of viable dark matter scenarios include: velocity-dependent (“p-wave”) annihilations, freeze-in dark matter (Hall et al, 2009; Elahi, Kolda, & Unwin, 2014) (for limits see Ginés, Witte, & Mena (2022); Kadota & Tashiro (2022); Chanda, Scholtz, & Unwin (2022); Chanda, Mukherjee, & Unwin (2025)), and (non-thermal) gravitationally coupled dark matter.
@@ 4.3 Probability of Close Encounter
For example, the probability that a flyby at initial distance 30 pc from the star passes within 15 AU of the star is roughly $p\sim p_{r_{c}}\sim 10^{-12}$. Figure 5 shows the probability that a PBH at initial distance $d=20,25,30$ pc passes within a given distance of closest approach ($\alpha$) of a star.
@@ 5 Distributions of Exoplanet Orbits Following Encounters
The perturbation strength decreases exponentially with respect to increases in the flyby velocity until the threshold value $R_{v}\approx 10$, whereafter the perturbation strength is insensitive to further increases in velocity. For a Sun-Jupiter system, an intruder of mass $M_{I}=10^{-2}M_{\odot}$ and velocity $|v_{0}|\approx 130$ km/s can lead to percent-level changes in planetary eccentricities ($\Delta e\sim 0.01$).
@@
The final eccentricity is determined by $e_{f}=e_{0}+\Delta e$, with our stated assumption $e_{0}$ this implies $e_{f}=\Delta e$. Notably, for close encounters with $e_{f}>1$, the planet will become unbound from the parent star. An interesting consequence of this is that such close encounters could lead to an excess of free-floating planets above those expected due to stellar close encounters. Interestingly, the OGLE telescope (Mroz, et al., 2017) reported a tentative excess of microlensing events corresponding to Earth-mass bodies. There is also an apparent excess of free-floating planets in the Upper Scorpius young stellar association (Miret-Roig et al., 2022). The OGLE excess was interpreted as an abundance of PBH (Niikura, et al., 2019; Scholtz & Unwin, 2020), but it could also be induced by an excess in free-floating planets (potentially arising from close encounters between PBH and stellar systems). Upcoming observations by the Rubin Observatory and Roman Space Telescope should provide new insights into populations of sub-stellar mass free-floating objects, with significant discovery potential.
Finally, we next turn to part C of Section 4 and examine how PBH flybys may alter the statistical distribution of exoplanet orbits. For each set of initial orbital parameters, we implement a Monte Carlo sampling technique to randomly choose the parameters of the flyby. For star-planet systems, the impact parameter $b$ of the flyby dictates the distance of closest approach and is drawn from a random distribution such that $b_{i}=r_{i}\times\sqrt{U[0,1]}$, where $r_{i}$ is the radius of critical interaction distance, and $U[0,1]$ is a random variable from a uniform distribution between 0 and 1. For star-Jupiter systems $r_{J}=15~{\rm AU}$, whereas for star-Neptune systems $r_{N}=90$ AU. The angular parameters $\Omega$, $\varpi$, and inclination $i$ of the flyby are all uniformly drawn from $\pi,\pi]$. Moreover, each initial setup is ran for $10$ samples, with each run independently sampling the angular orbital parameters.
Our findings are presented in Figure 9, which shows the distribution of changes to the eccentricity for Jupiter and Neptune-like planets assuming a PBH intruder with mass $10^{-1}M_{\odot}$. Changes to Jupiter’s eccentricity are strongly peaked at small values, whereas Neptune has a broader distribution. We attribute these differences to the fact that Jupiter is twenty times more massive than Neptune.
Refer to caption Refer to caption
Figure 9: The distribution of the resulting eccentricities of Jupiter (left)/Neptune (right) planet-star systems after a close encounter with a flyby PBH of mass $10^{-1}M_{\odot}$ and velocity $200$ km/s (taking for Jupiter $b_{J}\in 15\sqrt{U[0,1]}$AU and for Neptune $b_{N}\in 90\sqrt{U[0,1]}$AU).
@@ 7 Concluding Remarks
In this work we have utilysed the public package Gala to quantify the expected the number of PBH flybys for a given abundance and have shown that even with a small abundance ($f_{\rm PBH}\sim 10^{-7}$) a significant number of stars would have undergone a close encounter with a PBH. We then implemented a Monte Carlo study using a modified version of the public code Rebound to quantify how significantly the orbits would be perturbed. We leave for future work the interesting and feasible prospect of multiple PBH encounters on a system, and the prospect of a stellar encounter followed by PBH encounters. We would also like to examine more complicated stellar systems beyond the simplified Jupiter-Sun systems studied here.
@@
While we do not expect to be able to use exoplanet observations to place constraints in the near future, this work outlines the general principles of how one might use a future precision catalogue of exoplanets to discover or constrain populations of PBH. Finally, we note that if one could identify a class of clean planet-star systems that always formed with well-defined orbital parameters, these could provide an ideal testing ground for such dynamical constraints on PBH. Intriguingly, hot Jupiters seem to typically be found in single-planet systems (Wright et al., 2009; Steffen & Agol, 2005; Konacki et al., 2012), although exceptions have been observed; for instance, WASP-47b (Becker et al., 2015). Another potential route towards discriminatory power could be to look for objects with inclined orbits; however, even in our own solar system, we do not fully understand the origins of planetary orbits away from their ecliptic plane (see e.g. Batygin et al. (2019)). We leave these directions for future work.
@@ Acknowledgements
JU is supported by NSF grant PHY-2209998 and thanks the Berkeley Center for Theoretical Physics, Queen’s College, Oxford, and Rudolf Peierls Centre for Theoretical Physics, of the University of Oxford for their hospitality during this work. LH thanks the MIT-PRIMES program. We are grateful to Hanno Rein and Jakub Scholtz for helpful discussions and the anonymous referee for their useful remarks.
@@ Appendix B Multiple PBH Encounters
[…]
@@
[…]